FEL Upgrade Project Weekly Brief
June 18-22, 2001

This week’s FEL operations were devoted to the laser materials damage studies under contract to the DOD Joint Technology Office.  We began a large series exposures of various materials to FEL light at different irradiance levels from 1-10 kW/cm2.

Material generated at the high power FEL workshop and JTO working group meetings held earlier in the month were reviewed and condensed for us by the ONR Program Manager for presentation to the JTO next Tuesday.

Preparations were made for next week’s meeting of SURA’s Maritime Technical Advisory Committee (MTAC) which will be held on June 25-26 at Jefferson Lab.  MTAC is chaired by Vice Admiral Al Baciocco.  The committee has been and will continue to be very helpful to SURA and
Jefferson Lab as we plan our continuing partnership with the Navy.

Dr. Roger Mcginnis, the Navy representative to the JTO spent two days at the lab this week reviewing the first week of JTO sponsored work at the FEL.

A “surprise” safety inspection of the FEL facility was held today by representatives of the DOE Site Office, DOE’s Area Office in Oak Ridge and OSHA.  No violations were found and good marks were given for our training and procedures.  A useful recommendation was made for following
up on the disposition of chemicals and supplies that are brought to the FEL facility by users.

WBS 3 (Beam Physics)
We are exploring why the dumped beam is lower in energy than the injected energy (it’s because we power recover, not energy recover).  In the process, it became apparent that the power (not energy) recovery process utilizes only about half the available longitudinal aperture in the linac; a rebaseline of the process to energy recover the central energy of the distribution will require use of additional RF forward power, but should allow recovery of much larger momentum spread at the dump without loss.  The FEL thus may reasonably be expected to produce something like the 4 kW anticipated power limit on the mirrors.

WBS 4 (Injector)
We welded the mid-voltage rings for the ceramic stack and leak-checked them.  A small leak was found that should be easily repairable.

Gun HVPS - The upgraded unit was received.  It will be installed next year.  A new Gun Tank and Transmission Line elbow are being designed now.

Received prototype vacuum vessel and spaceframe. Acceptance inspection in progress.  Received reworked spaceframe tooling. Required repair/rework of aluminum welds.

WBS 6 (RF)
Zone 3 - The circulators for this zone were staged.  They will be installed within a few weeks.

Zone 4 - The 2.5 watt amplifiers were replaced with new 5 watt versions.  The new amplifiers will be tested when RF Control modules are available.  The missing circulator has been staged, but not yet

Injector HVPS - RFQ's for the transformers and inductors to upgrade the HVPS for the Quarter were sent to 7 prospective vendors.  Quotes are expected by next Friday.  The vendor for the 100 kW Klystrons is still unsure of his leakage problem.  He is remaking the parts for additional

WBS 8 (Instrumentation)
User assistance has been a priority this week as our new guests from the Naval Research Laboratory and Lincoln Lab begin their experiments with Michelle.

The installed Optical Beam Position Monitors appear to be working and will provide a useful tool for steering and monitoring beam position. The two that were calibrated using the drive laser are installed in Can 6 (Optics Control) and Can 2(downstairs in the beam transport line).  More testing is planned as a more stable running mode is achieved.

A User Power Supply was installed in Lab 1 for use by the folks working there.  It is providing a +5vdc trigger signal for their computer when the shutter leaves the shut position.  The computers they brought with them have been provided IP addresses so they can communicate with each other, but do not affect lab security, or the LAN.  An additional power supply is currently under construction.

Video was re-connected to the conference room television so the observers can watch from there as the control room was getting rather crowded.  Other user support was provided by operators on shift: James-2 days and one evening, Deborah-1 day and one evening.

The servo motor box has been worked, but is not yet complete.

Color plots of the VME Timing/Delay Circuit board have been completed and are awaiting review.  Components are on order to support assembly once the board is manufactured.

The assembly contract for the Ion Pump Power Supplies was awarded to Assembly and Design, Inc. in Richmond and parts were delivered Thursday.  We hope to see the early results back soon.

The assembly drawing for the controller board of the Ion Pump Power Supply was signed and submitted to document control prior to parts being sent to Assembly and Design, Inc.

Drawings at EECAD:  Molectron Quad Buffer Amplifier, Fabrication, Schematic and Assembly; Analog/Differential Driver Board Schematic; Pyro-electric Detector Buffer Amplifier, Schematic, Fabrication and Assembly; 32 Channel Sample/Hold Board (VME) Schematic, Assembly and
Fabrication; Ion Pump Power Supply Controller Board Schematic; Beam Viewer Chassis Test Box Schematic, Assembly and Fabrication; LSS PLC Test Box Schematic, Assembly and Fabrication; User Shutter Control Box Wiring Diagram; MPS System Diagram; Absolute Value/Sample and Hold Board Schematic, Assembly and Fabrication; LSS Wiring Diagram - Revision.

WBS 9 (Transport)
Shielded Bellows-- All components for the shielded bellows are in-house to begin fabrication.

Vacuum Chambers -- The interface document for all the vacuum chambers is ready for review by George Biallas. The next step is to refine the geometry to accommodate electron beam, photon beam, and magnet envelopes.

Assembly drawings continue for region 4F(Backleg) and 1F(Warm Region).

1. QX- Coil procurement-awarded last week, made initial contact with the vendor.
        Core procurement- in process of reviewing proposals.
        Pole chamfer tests- completed all machining that can be done on the prototype. Results have
        no converged to an acceptable solution.  Further review is needed.
2. Wiggler measurement stand-drawings in check
3. Quad measurement stand- drawings in progress
4. QT- no progress
5. SF-no progress

WBS 11 (Optics)
Most of the optics physics team (Michelle and Joe) was devoted to the execution of the JTO experiments this week.  With regard to the Upgrade:

    -detailing began on the internal mounting hardware in the optical cavity chambers
    -we received the second deformable mirror mount from the vendor; next step is to begin
      qualification tests using the Lab 2 laser
    -we reviewed drawings for the 2 inch deformable mirror mount and sent them on for continuing

The first two days of the week were used to set-up and qualify the sample transport system and other related hardware, which is being used to cycle the large number of samples for the JTO laser damage studies.  We were pleased to have a team from NRL involved with the experiments who has set-up a diagnostic for measuring plasma plumes from the irradiated samples and a team from Lincoln Lab who are involved in the sample analysis and modelling.  By 17:00 on Friday we completed 15
sample suites (with 6 samples per suite).  Sample effects are being measured on metal, ceramic and composite samples at irradiances between 1-10 kW/cm2.  Preliminary analysis shows sample melting on the stainless steel samples with the FEL beam where we would expect no melting with a cw laser beam at the same power and wavelength.  These experiments will continue all next week.